OpenStudy (anonymous):

Please explain this simply for me! Medical x-rays, with a wavelength of about 10^(-10) m, can penetrate completely through your skin. a. Ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength about 1,000 times the wavelength of an x-ray. Find the wavelength of ultraviolet rays. Show your work

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

10^(-10)(1000) = [10^(-10)](10^3)

OpenStudy (anonymous):

so what did you do

OpenStudy (anonymous):

the problem tells you to multiply the x-ray wavelenght by 1000

OpenStudy (anonymous):

they, of course, threw alot of unessicary information in there

OpenStudy (anonymous):

ya so all i have to do is multiply 10^-10 by 1000

OpenStudy (anonymous):

yeah, use the property of multiplying exponents with alike bases, x^a*x^b = x^(a+b)

OpenStudy (anonymous):

when i did it i got 1e-7

OpenStudy (anonymous):

10^(-7)m

OpenStudy (anonymous):

m = meters

OpenStudy (anonymous):

could you help me with the second part

OpenStudy (anonymous):

b. The wavelengths of visible light are between 3.8 times 10^-7m and 7.6 times 10^-7m. Are these wavelengths longer or shorter than those of ultraviolet rays? (The ultraviolet rays are 10^-7)

OpenStudy (anonymous):

this is testing your intuitive understanding of exponents, specificially negative ones

OpenStudy (anonymous):

remember x^-1=1/x

OpenStudy (anonymous):

yes so can you just help me with how to do a negative exponents like this one 12^-4

OpenStudy (anonymous):

12^-4 = 12^(4)(-1) = 1/12^4

OpenStudy (anonymous):

oh ok i get it

OpenStudy (anonymous):

(they are longer)